# ArtPro: Self-Supervised Articulated Object Reconstruction with Adaptive Integration of Mobility Proposals

### v1

@@ 3.3 Adaptive Proposal Integration Optimization
We define a series of loss terms as the optimization objectives. The RGBD image loss $\mathcal{L}_{I}$ and one-sided chamfer distance loss $\mathcal{L}_{cd}$ encourage the transformed Gaussians to align with the underlying surface of the end state, in order to ensure the reconstruction fidelity and enable the following proposal integration by checking the spatial relation of Gaussian centers. The part contrastive loss $\mathcal{L}_{pc}$ encourages the assignment of each Gaussian to a single dominant part by suppressing the probabilities of all non-maximum part proposals. The local smoothness loss $\mathcal{L}_{ls}$ enforces spatial smoothness on the static part probability field by encouraging similar values for neighboring Gaussian primitives. The regularization term $\mathcal{L}_{reg}$ encourages the compactness of each part by regularizing the part probabilities to align with a Gaussian distribution.
@@
$\mathcal{L}=\mathcal{L}_{I}+\lambda_{cd}\lambda\mathcal{L}_{cd}+\lambda_{pc}\mathcal{L}_{pc}+\lambda_{ls}\mathcal{L}_{ls}+\lambda_{reg}\mathcal{L}_{reg}.$ (7)
@@
$t_{m}\leftarrow t_{m}-(t_{m}\cdot a_{\text{col}})a_{\text{col}}.$ (8)
@@ 5 Conclusion
In this supplementary material, we first present the loss term formulations of Gaussian primitive optimization in Section A. Then, we present more two-part comparisons in Section B. In Section C and Section D, we provide more multi-part visualization and full initialization of our method on multi-part dataset. Finally, in Section E, we report the failure cases of our method.
@@ Appendix A Loss Terms of Gaussian Optimization
where $D_{v}(\cdot)$ denotes the rendered depth map of view $v$, $\mathcal{L}_{3dgs}$ is the RGB loss used in 3DGS [20], $\mathcal{G}$ is the Gaussian sets, $\{x_{m}\}$ is the Gaussian centers of all movable $\{\mathcal{G}_{m}\}$, and $\mathcal{P}^{1}$ is the point cloud of state $t=1$. We compute the loss $\mathcal{L}_{I}$ between the rendered images of transformed Gaussian $\mathcal{T}(\mathcal{G})$ and GT RGBD to optimize the appearance, and use $\mathcal{L}_{cd}$ to further improve the geometric quality.
@@
$\mathcal{L}=\mathcal{L}_{I}+\lambda_{cd}\lambda\mathcal{L}_{cd}+\lambda_{pc}\mathcal{L}_{pc}+\lambda_{ls}\mathcal{L}_{ls}+\lambda_{reg}\mathcal{L}_{reg}.$ (15)
@@
where $\mathcal{T}(\mathcal{G}_{m})$ is the union of all movable parts of $\mathcal{T}(\mathcal{G})$, $\mathcal{G}_{s}$ is the static part of $\mathcal{T}(\mathcal{G})$, $\alpha_{y}$ is the Gaussian opacity corresponding to center position $y$, and $\rm kNN$} is the set of $k$ nearest neighbors with $k=32$.
@@ Appendix B Additional Results on Two-Part Dataset
Figure 8 provides the full visualization results of ArtGS and Ours on this dataset. The results demonstrate that our method can robustly integrate the proposals and reconstruct both the part probability field and motion parameters. For incorrectly initialized mobility proposals, our optimization can also obtain correct motion estimation and the final reconstructions. For the real-world objects in this dataset, our method’s performance on the geometry reconstruction, especially in terms of CD metric, is limited by the quality of input depth maps. This is because our method tends to estimate the mobility parameters that align the transformed results closer to the depth map of state $t=1$. Large discrepancies between the input depth maps of states $t=0$ and $t=1$ lead to our results containing a small number of Gaussians originating from static part, such as the Real Storage in Figure 8 and Figure 13.
@@ Appendix D More Visual Analysis on Multi-Part Dataset
Appendix E Failure Cases
Although our approach achieves much more robust motion estimation and reconstruction of articulated objects than existing approaches, it still suffers from the common issues of the self-supervised 3DGS reconstruction framework. We show two representative failure cases in Figure 13. First, since our method lacks semantic part segmentation supervision for separating the adjacent movable parts, it sometimes causes the part boundary to encroach on neighboring part regions, such as the interior and the edge of the drawers in Figure 13. Second, the estimation of mobility parameters in our method primarily relies on the depth component of the RGBD loss $\mathcal{L}_{I}$ and the geometry-aware Chamfer distance loss $\mathcal{L}_{cd}$. Consequently, the accuracy of these estimations is highly dependent on the input depth map quality. In scenarios where the acquired depth maps contain significant inaccuracies, due to factors such as sensor noise, occlusions, or varying illumination conditions, the optimization process can be misled. As illustrated in Figure 13 (using real-world data from the PARIS dataset), this can cause our method to erroneously incorporate Gaussians from the static part into the estimated movable part. To achieve a lower loss value under these incorrect constraints, the optimization incorrectly transforms the movable part into the interior of the object. It is worth noting that in our real-world reconstruction experiments, the quality of the mobility results has been significantly improved by enhancing the input depth maps using the pre-trained Depth-Anything-V2 model [52] which effectively mitigates these issues.

### v2

@@ 3.3 Adaptive Proposal Integration Optimization
We define a series of loss terms as the optimization objectives. The RGBD image loss $\mathcal{L}_{I}$ of end state and one-sided chamfer distance loss $\mathcal{L}_{cd}$ encourage the transformed Gaussians to align with the underlying surface of the end state, in order to ensure the reconstruction fidelity and enable the following proposal integration by checking the spatial relation of Gaussian centers. The part contrastive loss $\mathcal{L}_{pc}$ encourages the assignment of each Gaussian to a single dominant part by suppressing the probabilities of all non-maximum part proposals. The local smoothness loss $\mathcal{L}_{ls}$ enforces spatial smoothness on the static part probability field by encouraging similar values for neighboring Gaussian primitives. The regularization term $\mathcal{L}_{reg}$ encourages the compactness of each part by regularizing the part probabilities to align with a Gaussian distribution.
@@
$\mathcal{L}=\mathcal{L}_{I}+\lambda_{cd}\mathcal{L}_{cd}+\lambda_{pc}\mathcal{L}_{pc}+\lambda_{ls}\mathcal{L}_{ls}+\lambda_{reg}\mathcal{L}_{reg}.$ (7)
@@
$t_{m}\leftarrow t_{m}-(t_{m}\cdot a_{\text{col}})\cdot a_{\text{col}}.$ (8)
@@ 5 Conclusion
6 Acknowledgement
This work is supported by the National Natural Science Foundation of China (62302269), the Excellent Young Scientists Fund Program (Overseas) of Shandong Province (No.2023HWYQ-034), the Joint Funds of the National Natural Science Foundation of China (U23A20312), a grant from the Chongqing Natural Science Foundation (CSTB2024NSCQ-MSX1026), the Natural Science Foundation of Shandong Province (No.ZR2023QF077), Shandong Province Key Research and Development Program (2024TSGC0051). We also acknowledge the support from the grant by Lejoin Intelligence (Shenzhen) Co.,Ltd.
@@ 6 Acknowledgement
In this supplementary material, we first present the loss term formulations of Gaussian primitive optimization in Section A. Then, we present more two-part comparisons in Section B. In Section C and Section D, we provide more multi-part visualization and full initialization of our method on multi-part dataset. In Section E and Section F, we report additional ablation studies and computational costs. Finally, in Section G, we report the failure cases of our method.
@@ Appendix A Loss Terms of Gaussian Optimization
where $D_{v}(\cdot)$ denotes the rendered depth map of view $v$, $\mathcal{L}_{3dgs}$ is the RGB loss used in 3DGS [20], $\mathcal{G}$ is the Gaussian sets, $\{x_{m}\}$ is the Gaussian centers of all movable $\{\mathcal{G}_{m}\}$, and $\mathcal{P}^{1}$ is the point cloud of end state. We compute the loss $\mathcal{L}_{I}$ between the rendered images of transformed Gaussian $\mathcal{T}(\mathcal{G})$ and GT RGBD of end state to optimize the appearance, and use $\mathcal{L}_{cd}$ to further improve the geometric quality.
@@
$\mathcal{L}=\mathcal{L}_{I}+\lambda_{cd}\mathcal{L}_{cd}+\lambda_{pc}\mathcal{L}_{pc}+\lambda_{ls}\mathcal{L}_{ls}+\lambda_{reg}\mathcal{L}_{reg}.$ (15)
@@
where $\mathcal{T}(\mathcal{G}_{m})$ is the union of all movable parts of $\mathcal{T}(\mathcal{G})$, $\mathcal{G}_{s}$ is the static part of $\mathcal{T}(\mathcal{G})$, $\hat{\alpha}_{y}$ is the Gaussian opacity corresponding to center position $y$, and $\rm kNN$} is the set of $k$ nearest neighbors with $k=32$.
@@ Appendix B Additional Results on Two-Part Dataset
Figure 8 provides the full visualization results of ArtGS and Ours on this dataset. The results demonstrate that our method can robustly integrate the proposals and reconstruct both the part probability field and motion parameters. For incorrectly initialized mobility proposals, our optimization can also obtain correct motion estimation and the final reconstructions. For the real-world objects in this dataset, our method’s performance on the geometry reconstruction, especially in terms of CD metric, is limited by the quality of input depth maps. This is because our method tends to estimate the mobility parameters that align the transformed results closer to the depth map of end state ($t=1$). Large discrepancies between the input depth maps of states $t=0$ and $t=1$ lead to our results containing a small number of Gaussians originating from static part, such as the Real Storage in Figure 8 and Figure 13.
@@ Appendix D More Visual Analysis on Multi-Part Dataset
We further report the visual quality metrics on our dataset in Table 5. Our method achieves better PSNR, SSIM, and LPIPS scores compared to ArtGS, while producing more accurate articulation.
Table 5: Visual quality on our dataset.
Method PSNR SSIM LPIPS
ArtGS 36.10 0.980 0.037
Ours 48.02 0.998 0.003
Appendix E More Ablation Study Results
To strengthen the validation of careful initialization, we report the results of “w/o MoI” (no principal axes for initialization) and “w/o MoI&\&PI” (no principal axes for initialization and pruning) in Table 6. We also add the results of disabling the progressive merging after over-segmentation (denoted as w/o PM) in Table 6.
Table 6: Additional ablation study on our dataset.
Case CD-s $\overline{\text{CD}}$-m CD-w
w/o PM 0.85 4.26 0.68
w/o MoI 0.89 5.62 0.70
w/o MoI&PI 0.87 5.81 0.71
Full 0.84 3.63 0.65
[…]
Method Time (min) GPU Memory (GB)
[…]
Appendix F Computation Overhead
We control the computational cost of ArtPro through two key strategies. First, we use PartField [29] to produce well-structured over-segmentation proposals rather than fragmented patches, and tend to merge the most relevant proposals in the first optimization cycle, which limits the total number of proposals and cycles. Second, the optimization of each cycle is monitored and stops early upon convergence.
Table 7 reports the average computation time and GPU memory usage across all objects in our dataset. Our method achieves a practical balance between computational cost and reconstruction quality. Although ArtGS is faster, it cannot robustly reconstruct complex multi-part objects, as shown in Tables 1–2 of the mainpage. DTA, on the other hand, requires substantially more time and memory while also failing on many multi-part cases.
@@ Appendix F Computation Overhead
Figure 14: Influence of noisy depth and mask.
Appendix G Failure Cases and Robustness Analysis
Robustness to noisy depth and masks. Fig. 14 shows the robustness of our method to noisy depth maps and imprecise segmentation masks. For local geometry errors (Fig. 14(a)) and imprecise mask boundaries near drawer edges (Fig. 14(c)), our method mitigates the incorrect estimation from specific views by integrating multi-view information, achieving accurate reconstructions. However, our method would fail when the captured depth maps and masks are highly fragmented and misaligned, as in the real-scanned data provided by PARIS [25]. This can be enhanced using prior-prompted depth predictor, e.g., PriorDA [45], as we did to obtain our real-scanned data (the printer and storage).
Part boundary sensitivity. Although our approach achieves much more robust motion estimation and reconstruction of articulated objects than existing approaches, it still suffers from the common issues of the self-supervised 3DGS reconstruction framework. We show two representative failure cases in Figure 13. First, since our method lacks semantic part segmentation supervision for separating the adjacent movable parts, it sometimes causes the part boundary to encroach on neighboring part regions, such as the interior and the edge of the drawers in Figure 13.
Sensitivity to depth map quality. Second, the estimation of mobility parameters in our method primarily relies on the depth component of the RGBD loss $\mathcal{L}_{I}$ and the geometry-aware Chamfer distance loss $\mathcal{L}_{cd}$. Consequently, the accuracy of these estimations is highly dependent on the input depth map quality. In scenarios where the acquired depth maps contain significant inaccuracies, due to factors such as sensor noise, occlusions, or varying illumination conditions, the optimization process can be misled. As illustrated in Figure 13 (using real-world data from the PARIS dataset), this can cause our method to erroneously incorporate Gaussians from the static part into the estimated movable part. To achieve a lower loss value under these incorrect constraints, the optimization incorrectly transforms the movable part into the interior of the object. It is worth noting that in our real-world reconstruction experiments, the quality of the mobility results has been significantly improved by enhancing the input depth maps using the pre-trained Depth-Anything-V2 model [53] which effectively mitigates these issues.
Multi-DOF joints and non-rigid parts. Our current formulation assumes that each movable part undergoes a single rigid transformation, either revolute or prismatic. This assumption does not hold for objects with multi-DOF joints such as ball joints or compound hinges, or parts exhibiting non-rigid deformations. In such cases, our motion parameterization cannot capture the full range of part motion, leading to inaccurate articulation estimation. Future work will explore potential solutions including extending the motion model to support composite transformations or integrating learned deformation fields for non-rigid components.